A MASS OF 4.00 KG, INITALLY At rest, explodes. A 1.50 kg piece moves at 4.00 m/s at 60.0 degrees north of west. A 0.50 kg piece moves at 8.00 m/s east.

The magnitude of the momentum of the third piece is _____

please show work.
Answer #1

By momentum conservation:-

Initial momentum= 0

Final momentum also be zero.

Momentum of 1st particle:-

P1= 1.5(-4cos(60)i+4sin(60)j)= -3i+5.196j kg-m/s

P2= 4i kg-m/s

Momentum of third particle=

P3= -(P1+P2) = i+5.196j kg-m/s

P3= 5.29 kg-m/s